# JWST imaging of the Pleiades: anisotropy of turbulence in the cold neutral medium

### v1

@@ 3.3 CIB residual
Figures 4 and 5 indicate that the power spectra include an isotropic emission component, which prevails at high values. We characterize this component, which we identify as an extragalactic residual left after data cleaning.
@@
Notes. (a): Amplitude of power-law at wavenumber pix. (b): Spectral index of the power-law. (c): Reduced of the fit.
@@
The spectral index of the isotropic component, about , coincides with the value measured for the clustering of faint extragalactic sources combined into a diffuse extragalactic background. The cosmic near-infrared background (CIB) was detected with IRAC Spitzer data. The measurements reported by Kashlinsky et al. (2012) for the IRAC m channel can be directly compared with the results of our JWST analysis. The spectral index of the isotropic component is consistent with the Spitzer CIB power spectra, and the amplitude is one order of magnitude smaller. We conclude that the isotropic component is a residual CIB, which remains after the masking of extragalactic sources666We note that the spectra in Fig. 6 do not exhibit flattening at high , which could be interpreted as shot noise from faint sources at the JWST detection limit..
@@
Notes. (a): Amplitude of power-law at wavenumber pix. (b): Spectral index of the power-law. (c): Mask used to clean the data. F, L and H stand for fiducial, light and heavy masks (see Sect. 2.2). (d): Median brightness at m of diffuse emission computed on the Field 1 - Field 2 difference maps.
@@ 4 Power spectra of PAH emission
In this section, we derive power spectra of the Pleiades PAH emission using JWST data. To obtain these spectra, we need to subtract the CIB residual. This is achieved through three distinct methods, each based on different hypotheses. In Sect. 4.1, we assume that the CIB residual is the same in our two fields, both close to and distant from Merope, and we use the latter for subtraction. In Sect. 4.2, we use the anisotropy of the Pleiades emission to subtract the isotropic CIB residual. Sect. 4.3 presents fits of the Field 1 spectra using two power laws to represent the PAH contribution and the CIB residual.
@@
Notes.(a): Parameters of the power-law fit to for in Eq. 8. (b): Masks used to compute power spectra. F, L and H stand for fiducial, light and heavy.
4.1 Power spectra from spatial difference
For each spectrum obtained for Cuts 1 and 2 of Field 1, we subtract the mean of the Field 2 power spectra for Cuts 1 and 2. The difference subtracts the CIB residual and a potential contribution from the Galactic background to the PAH spectra. This approach does not make use of the anisotropy fit, which restricts the range of bins that can be measured with relevant accuracy. The differences in spectra are presented in Fig. 7. The dashed lines represent power-law fits:
@@ 4.1 Power spectra from spatial difference
The fit parameters, the amplitude and the spectral index , are listed in Table 4. The table also includes the mean of the sky emission measured after source masking and OFF subtraction. Figure 7 shows the spectra, and Table 4 lists the fit results, for the three masks. The light and heavy masks are used to assess systematic uncertainties. The data points in Fig. 7 show a noticeable steepening of the spectra at high for the heavy mask. The difference in the specral index is slightly larger than the statistical error. This may be due to the lack of small-scale structure in the multiscale background used to fill in the masked areas.
In this work, we assume that on the spatial scales probed by the Pleiades JWST data, PAH emission traces the gas column density. For scales smaller than the line-of-sight depth of the emitting medium, the spectral index of the column density power spectrum is identical to that of the three-dimensional density field (Cho and Lazarian, 2003; Brunt et al., 2010). We therefore proceed under the assumption that the PAH power spectrum is a proxy for the gas density power spectrum.
The spectral index that we measure matches the reference spectral index for MHD turbulence (Iroshnikov, 1964; Kraichnan, 1965). It is also consistent with the index inferred by Armstrong et al. (1995) for electron density fluctuations in the local interstellar medium by analyzing pulsar scintillation measurements on scales below mpc.
To compare with previous measurements of dust power spectra of the diffuse ISM, we refer to the spectra measured by Gibson (2007) in the Pleiades and by Miville-Deschênes et al. (2016) for an independent field at high Galactic latitude. These spectra, which were obtained by analyzing images of dust scattered light, are the closest observational match to our analysis in terms of angular resolution. For far-infrared dust emission, we also consider results derived Herschel observations by Miville-Deschênes et al. (2010) and Auclair et al. (2024) with angular resolutions of 30 and , respectively. The spectra reported in these four dust studies with spectral indices ranging from -2.7 to -2.95 are shallower than the two shown in Figure 7 .
@@ 4.2 Power spectra from anisotropy
In our second approach, we assume that the residual CIB is isotropic and obtain power spectra for each of the JWST images taking the difference . These differences are shown in Fig. 8. To highlight the difference in slope between the Field 1 and 2 images, the spectra are scaled by . The dashed-lines represent power-law fits to the spectra:
The fit parameters and for each of the images are listed in Table 5. We also list values for the light and heavy masks to assess systematic errors. The differences with the values for the fiducial mask are minor. It is satisfying to obtain values for the power law indices of the spectra, which are consistent with those obtained for the Field 1 cuts using the spatial difference (see Table 4).
4.3 Two components fit
@@
We perform a Monte Carlo Markov chain analysis (MCMC) to sample the joint posterior distribution of the model parameters. We utilize the EMCEE python package as described by Foreman-Mackey et al. (2013) in accordance with their recommendations. We use a Gaussian likelihood. The priors force the three amplitudes , and to be positive, and the spectral indices and to be negative. In addition, we set a Gaussian prior in the spectral index with a dispersion of 0.2 (). This prior follows from the fit of for Field 2 images illustrated in Fig. 6. As this prior makes use of Field 2 data, we run the MCMC analysis only for the two Field 1 images.
@@
The model parameters with their error bars are listed in Table 6. The table lists the median values of the posterior distribution and error bars computed from the 16 and 84 percentiles. As shown in Fig. 9 for Field 1 Cut 1, the model provides a good fit for both and . The amplitude of the CIB residual is consistent with that measured for parallel spectra of Field 2, within error bars (see Table 3). The values of also match the spectral indices derived in Sects. 4.1 and 4.2. It is satisfying to validate these values using a distinct approach in which the CIB residual is included in the data model with a power-law spectrum, rather than being subtracted by a difference. We note that the degree of anisotropy, measured by the ratios in Table 6, differs by a factor 2 between Cuts 1 and 2 .
@@
Figure 10 shows the joint posterior distributions of and , derived from the MCMC fits of Field 1 Cuts 1 and 2. The difference between the two spectral indices depends on and, more generally, on the simplified model used to separate the PAH spectra from the CIB residual. In Figure 10, the median value of is slightly negative, but we consider the deviation from zero to be not statistically significant. We conclude that the two spectral indices are identical within a measurement error bar of 0.2 ().
@@
As Lithwick and Goldreich (2001)777See Sect. 3 in this publication pointed out in the context of their interpretation of radio scintillation observations, the observational effects of the parallel direction are washed out due to variations in the local orientation of anisotropy along the line of sight and across the sky. In our observations, the PAH emission is so highly anisotropic that modes perpendicular to the local orientation of anisotropy dominate the Fourier spectra in all projected orientations onto the plane of the sky. This statement entails three key facts. First, the ratios in Table 6 are lower limits to the intrinsic anisotropy of the PAH emission. These limits depend on the dispersion of the local orientation of the anisotropy within the emitting medium. Second, the spectral index in Eq. 9 does not represent the spectral index of the energy cascade in the direction parallel to the magnetic field. Third, the finding that is consistent with 0 does not exclude the possibility that the spectral indices of the turbulent energy cascade for perpendicular and parallel modes differ.
@@
Notes. Parameters of the joint fits of and for Field 1 images Cuts 1 and 2, using data model in Eq. 9 with pix-1.
@@ 5 Relation to interstellar turbulence
In this section, we present three complementary perspectives that relate the JWST Pleiades observations to interstellar turbulence. Together, they provide a framework that we hope will motivate theoretical studies aimed at interpreting the data.
@@ 5.1 Dynamical interaction between the Pleiades stars and the surrounding ISM
Projection effects make it difficult to infer the distance at which the interaction with the stars occurs. Modeling of scattered light from dust indicates that the distance between matter and the brightest Pleiades stars ranges from a few tenths to 1 parsec. Specifically, it is a few tenths of a parsec near Merope (Gibson and Nordsieck, 2003b), a value that is consistent with our estimate of the radiation field intensity at the position of the Field 1 images (see Sect. C). For Merope’s luminosity, a value of G in the range 10–20 corresponds to a distance from the star between 0.3 and 0.5 pc.
@@ 5.2 CNM structure on the scales of the Pleiades JWST observations
To support this statement, we compute the eddy turnover timescale for a power law scaling of the velocity dispersion:
where is the dispersion of turbulent velocity on the scale and is the scale of energy injection. We use the size of the Pleiades nebula for , and the dispersion of velocities among gas tracers reported by Ritchey et al. (2006) as an observational estimate of .
We assume that magnetic fields are dynamically important on the scale of the Pleiades nebula. We therefore use the canonical scaling of the velocity dispersion for MHD turbulence (Iroshnikov, 1964; Kraichnan, 1965). For this index, the velocity dispersion is subsonic on the scales (mpc) statistically sampled by our JWST observations. The eddy turnover timescale is as follows:
We compare with the gas cooling time, which we computed for a reference pressure in the local ISM using the analytical expression of the cooling function in Jennings and Li (2021). For mpc, is at least two orders of magnitude smaller than the cooling timescale of the warm neutral medium. For mpc, is comparable to the cooling time of the CNM, and shorter for lower values of . It is also relevant to estimate the Field length . Thermal instability is suppressed by conduction on scales smaller than .
@@ 5.3 MHD subsonic turbulence
The anisotropy power ratios relative to this mean orientation, 90 and 40 for Fields 1 and 2 (Table 6), are lower bounds to the local anisotropy, due to the dispersion in the orientation of the striations (Sect. 3.2). This dispersion – as determined by the widths of the angular profiles of the 2 dimensional power spectra – is significant (see Table 2).
The anisotropy of is observed to be scale independent within a small uncertainty set by the separation between the PAH emission and the CIB residual. We stress that for a large power anisotropy, we expect perpendicular modes to be dominant in all orientations projected onto the plane of the sky (Sect. 4.3)
@@
The first interpretation was quantified by Beattie and Federrath (2020) using numerical simulations. More research involving dedicated simulations is required to verify whether this interpretation can reproduce the main observational results. The second is based on the Goldreich and Sridhar (1995) model of subsonic MHD turbulence. In this framework, compressible modes are sheared and passively mixed by Alfvén waves (Lithwick and Goldreich, 2001). This interpretation is attractive because it links the JWST observations to the anisotropic nature of the MHD turbulence energy cascade (Boldyrev, 2006; Schekochihin, 2022).
@@ 6 Conclusion
We build maps of diffuse PAH emission, separating it from sources, stars, and galaxies, as well as the extended spikes of bright stars. The stars were subtracted using an empirical template of the JWST PSF. Galaxies are identified and masked using multiresolution image filtering to interpolate the structure of diffuse emission.
We present a power spectrum analysis of the diffuse emission in four images, two of which are close to the Merope star and two of which are more distant. The spectra exhibit flattening at high wavenumbers, which we identify as a CIB residual. The spectral index of this component, -1.2, is similar to that measured for galaxy clustering in CIB observations at near- and far-infrared wavelengths.
The amplitude of PAH power spectra are highly anisotropic on all scales. We characterize them in three independent ways. (1) We subtract the CIB residual by taking the difference of the spectra computed for the JWST images near and far from Merope. (2) We differentiate anisotropic PAH emission from the isotropic CIB residual by modeling the 2D angular distribution of Fourier power. (3) We fit the power spectra in the directions parallel and perpendicular to the anisotropy orientation as the sum of two power laws, representing the contributions from PAHs and the CIB. The three methods yield consistent results.
The PAH power spectra are measured to a spatial resolution of 0.2 mpc (40 au). They are well fitted with single power laws, with no sign of a break that could be interpreted as a signature of a characteristic scale associated with turbulence dissipation. The power laws indices are -3.5 in the two fields near Merope and -3 in the more distant areas. The magnetic field orientation inferred from dust polarization data from Planck is aligned with the PAH anisotropy. The power anisotropy does not change with scale.
These findings are discussed in the context of interstellar turbulence, which may be driven by the dynamical interaction between the Pleiades stars and the surrounding ISM. The observations offer a new perspective for comparing observations and theoretical models because they reveal physical scales at which CNM turbulence is decoupled from the thermal instability. Progress in interpreting the data within this framework requires follow-up theoretical work including numerical simulations to link the PAH power spectrum to the three-dimensional, anisotropic structure of the magnetized ISM.
@@ Acknowledgements.
GV, AAC and NDR thank the STScI for DDRF Summer funding to support their research. We thank the referee for a detailed report, which helped us to clarify the discussion of the results of the data analysis in relation to interstellar turbulence.References
@@
- A Full-sky, High-resolution Atlas of Galactic 12 m Dust Emission with WISE. ApJ 781 (1), pp. 5. External Links: Document, 1312.0947, ADS entry Cited by: Table 7, Appendix B, Appendix C.

### v2

@@ 3.3 Cosmic near-infrared background residual
Figures 4 and 5 indicate that the power spectra include an isotropic emission component, which prevails at high values. We characterized this component, which we identified as an extragalactic residual left after data cleaning.
@@
The spectral index of the isotropic component, about , coincides with the value measured for the clustering of faint extragalactic sources combined into a diffuse extragalactic background. The cosmic near-infrared background (CIB) was detected with Infrared Array Camera (IRAC) Spitzer data. The measurements reported by Kashlinsky et al. (2012) for the IRAC m channel can be directly compared with the results of our JWST analysis. The spectral index of the isotropic component is consistent with the Spitzer CIB power spectra, and the amplitude is one order of magnitude smaller. We conclude that the isotropic component is a residual CIB, which remains after the masking of extragalactic sources999We note that the spectra in Fig. 6 do not exhibit flattening at high , which could be interpreted as shot noise from faint sources at the JWST detection limit..
@@ 4 Power spectra of PAH emission
In this section we derive power spectra of the Pleiades PAH emission using JWST data. To obtain these spectra, we needed to subtract the CIB residual. This was achieved through three distinct methods, each based on different hypotheses. In Sect. 4.1 we assume that the CIB residual is the same in our two fields, both close to and far from Merope, and we used the latter for subtraction. In Sect. 4.2 we use the anisotropy of the Pleiades emission to subtract the isotropic CIB residual. Section 4.3 presents fits of the Field 1 spectra using two power laws to represent the PAH contribution and the CIB residual.
@@
4.1 Power spectra from the spatial difference
For each spectrum obtained for Cuts 1 and 2 of Field 1, we subtracted the mean of the Field 2 power spectra for Cuts 1 and 2. The difference subtracts the CIB residual and a potential contribution from the Galactic background to the PAH spectra. This approach does not make use of the anisotropy fit, which restricts the range of bins that can be measured with relevant accuracy. The differences in spectra are presented in Fig. 7. The dashed lines represent power-law fits:
@@ 4.1 Power spectra from the spatial difference
The fit parameters, the amplitude and the spectral index , are listed in Table 4. The table also includes the mean of the sky emission measured after source masking and OFF subtraction. Figure 7 shows the spectra, and Table 4 lists the fit results, for the three masks. The light and heavy masks are used to assess systematic uncertainties. The data points in Fig. 7 show a noticeable steepening of the spectra at high for the heavy mask. The difference in the spectral index is slightly larger than the statistical error. This may be due to the lack of small-scale structure in the multi-scale background used to fill in the masked areas.
In this work, we assumed that on the spatial scales probed by the Pleiades JWST data, PAH emission traces the gas column density. For scales smaller than the line-of-sight depth of the emitting medium, the spectral index of the column density power spectrum is identical to that of the 3D density field (Cho and Lazarian, 2003; Brunt et al., 2010). We therefore proceed under the assumption that the PAH power spectrum is a proxy for the gas density power spectrum.
The spectral index that we measure matches the reference spectral index for MHD turbulence (Iroshnikov, 1964; Kraichnan, 1965). It is also consistent with the index inferred by Armstrong et al. (1995) for electron density fluctuations in the local ISM by analyzing pulsar scintillation measurements on scales below mpc.
To compare with previous measurements of dust power spectra of the diffuse ISM, we refer to the spectra measured by Gibson (2007) in the Pleiades and by Miville-Deschênes et al. (2016) for an independent field at high Galactic latitude. These spectra, which were obtained by analyzing images of dust scattered light, are the closest observational match to our analysis in terms of angular resolution. For far-infrared dust emission, we also consider results derived from Herschel observations by Miville-Deschênes et al. (2010) and Auclair et al. (2024) with angular resolutions of 30 and , respectively. The spectra reported in these four dust studies with spectral indices ranging from -2.7 to -2.95 are shallower than the two shown in Fig. 7 .
@@ 4.2 Power spectra from anisotropy
In our second approach, we assume that the residual CIB is isotropic and obtain power spectra for each of the JWST images by taking the difference . These differences are shown in Fig. 8. To highlight the difference in slope between the Field 1 and 2 images, the spectra are scaled by . The dashed lines represent power-law fits to the spectra:
The fit parameters and for each of the images are listed in Table 5. We also list values for the light and heavy masks to assess systematic errors. The differences with the values for the fiducial mask are minor. It is satisfying to obtain values for the power-law indices of the spectra, which are consistent with those obtained for the Field 1 cuts using the spatial difference (see Table 4).
4.3 Two-component fit
@@
We performed a Monte Carlo Markov chain (MCMC) analysis to sample the joint posterior distribution of the model parameters. We utilized the EMCEE python package as described by Foreman-Mackey et al. (2013) in accordance with their recommendations. We used a Gaussian likelihood. The priors force the three amplitudes , , and to be positive, and the spectral indices and to be negative. In addition, we set a Gaussian prior in the spectral index with a dispersion of 0.2 (). This prior follows from the fit of for Field 2 images illustrated in Fig. 6. As this prior makes use of Field 2 data, we ran the MCMC analysis only for the two Field 1 images.
@@
The model parameters with their error bars are listed in Table 6. The table lists the median values of the posterior distribution and error bars computed from the 16 and 84 percentiles. As shown in Fig. 9 for Field 1 Cut 1, the model provides a good fit for both and . The amplitude of the CIB residual is consistent with that measured for parallel spectra of Field 2, within error bars (see Table 3). The values of also match the spectral indices derived in Sects. 4.1 and 4.2. It is satisfying to validate these values using a distinct approach in which the CIB residual is included in the data model with a power-law spectrum, rather than being subtracted by a difference. We note that the degree of anisotropy, measured by the ratios in Table 6, differs by a factor of 2 between Cuts 1 and 2 .
@@
Figure 10 shows the joint posterior distributions of and , derived from the MCMC fits of Field 1 Cuts 1 and 2. The difference between the two spectral indices depends on and, more generally, on the simplified model used to separate the PAH spectra from the CIB residual. In Fig. 10, the median value of is slightly negative, but we consider the deviation from zero to be not statistically significant. We conclude that the two spectral indices are identical within a measurement error bar of 0.2 ().
@@
As Lithwick and Goldreich (2001)121212See Sect. 3 in this publication pointed out in the context of their interpretation of radio scintillation observations, the observational effects of the parallel direction are washed out due to variations in the local orientation of anisotropy along the line of sight and across the sky. In our observations, the PAH emission is so highly anisotropic that modes perpendicular to the local orientation of anisotropy dominate the Fourier spectra in all projected orientations onto the plane of the sky. This statement entails three key facts. First, the ratios in Table 6 are lower limits to the intrinsic anisotropy of the PAH emission. These limits depend on the dispersion of the local orientation of the anisotropy within the emitting medium. Second, the spectral index in Eq. 9 does not represent the spectral index of the energy cascade in the direction parallel to the magnetic field. Third, the finding that is consistent with 0 does not exclude the possibility that the spectral indices of the turbulent energy cascade for perpendicular and parallel modes differ.
@@ 5 Relation to interstellar turbulence
In this section we present three complementary perspectives that relate the JWST Pleiades observations to interstellar turbulence. Together, they provide a framework that we hope will motivate theoretical studies aimed at interpreting the data.
@@ 5.1 Dynamical interaction between the Pleiades stars and the surrounding ISM
Projection effects make it difficult to infer the distance at which the interaction with the stars occurs. Modeling of scattered light from dust indicates that the distance between matter and the brightest Pleiades stars ranges from a few tenths to 1 parsec. Specifically, it is a few tenths of a parsec near Merope (Gibson and Nordsieck, 2003b), a value that is consistent with our estimate of the radiation field intensity at the position of the Field 1 images (see Appendix C). For Merope’s luminosity, a value of G in the range 10–20 corresponds to a distance from the star between 0.3 and 0.5 pc.
@@ 5.2 CNM structure on the scales of the Pleiades JWST observations
To support this statement, we computed the eddy turnover timescale for a power law scaling of the velocity dispersion:
where is the dispersion of turbulent velocity on the scale and is the scale of energy injection. We used the size of the Pleiades nebula for , and the dispersion of velocities among gas tracers reported by Ritchey et al. (2006) as an observational estimate of .
We assume that magnetic fields are dynamically important on the scale of the Pleiades nebula. We therefore used the canonical scaling of the velocity dispersion for MHD turbulence (Iroshnikov, 1964; Kraichnan, 1965). For this index, the velocity dispersion is subsonic on the scales (mpc) statistically sampled by our JWST observations. The eddy turnover timescale is as follows:
We compared with the gas cooling time, which we computed for a reference pressure in the local ISM using the analytical expression of the cooling function in Jennings and Li (2021). For mpc, is at least two orders of magnitude smaller than the cooling timescale of the warm neutral medium. For mpc, is comparable to the cooling time of the CNM, and shorter for lower values of . It is also relevant to estimate the Field length (). Thermal instability is suppressed by conduction on scales smaller than .
@@ 5.3 MHD subsonic turbulence
The anisotropy power ratios relative to this mean orientation, 90 and 40 for Fields 1 and 2 (Table 6), are lower bounds to the local anisotropy, due to the dispersion in the orientation of the striations (Sect. 3.2). This dispersion – as determined by the widths of the angular profiles of the 2D power spectra – is significant (see Table 2).
The anisotropy of is observed to be scale-independent within a small uncertainty set by the separation between the PAH emission and the CIB residual. We stress that for a large power anisotropy, we expect perpendicular modes to be dominant in all orientations projected onto the plane of the sky (Sect. 4.3)
@@
The first interpretation was introduced and quantified using numerical simulation in the context of molecular cloud observations by Tritsis and Tassis (2016) and Tritsis et al. (2018). It was also investigated by Beattie and Federrath (2020) using numerical simulations. More research involving dedicated simulations is required to verify whether this interpretation can reproduce the main observational results. The second is based on the Goldreich and Sridhar (1995) model of subsonic MHD turbulence. In this framework, compressible modes are sheared and passively mixed by Alfvén waves (Lithwick and Goldreich, 2001). This interpretation is attractive because it links the JWST observations to the anisotropic nature of the MHD turbulence energy cascade (Boldyrev, 2006; Schekochihin, 2022).
@@ 6 Conclusion
We built maps of diffuse PAH emission, separating it from sources, stars, and galaxies, as well as from the extended spikes of bright stars. The stars were subtracted using an empirical template of the JWST PSF. Galaxies were identified and masked using multi-resolution image filtering to interpolate the structure of diffuse emission.
We have presented a power spectrum analysis of the diffuse emission in four images, two of which are close to the Merope star and two of which are more distant. The spectra exhibit flattening at high wavenumbers, which we identify as a CIB residual. The spectral index of this component, -1.2, is similar to that measured for galaxy clustering in CIB observations at near- and far-infrared wavelengths.
The amplitude of PAH power spectra are highly anisotropic on all scales. We characterized them in three independent ways: (1) We subtracted the CIB residual by taking the difference of the spectra computed for the JWST images near and far from Merope. (2) We differentiated anisotropic PAH emission from the isotropic CIB residual by modeling the 2D angular distribution of Fourier power. (3) We fit the power spectra in the directions parallel and perpendicular to the anisotropy orientation as the sum of two power laws, representing the contributions from PAHs and the CIB. The three methods yield consistent results.
The PAH power spectra are measured to a spatial resolution of 0.2 mpc (40 au). They are well fitted with single power laws, with no sign of a break that could be interpreted as a signature of a characteristic scale associated with turbulence dissipation. The power-law indices are -3.5 in the two fields near Merope and -3 in the more distant areas. The magnetic field orientation inferred from dust polarization data from Planck is aligned with the PAH anisotropy. The power anisotropy does not change with scale.
These findings are discussed in the context of interstellar turbulence, which may be driven by the dynamical interaction between the Pleiades stars and the surrounding ISM. The observations offer a new perspective for comparing observations and theoretical models because they reveal physical scales at which CNM turbulence is decoupled from the thermal instability. Progress in interpreting the data within this framework requires follow-up theoretical work, including numerical simulations to link the PAH power spectrum to the 3D anisotropic structure of the magnetized ISM.
@@ Acknowledgements.
GV, AAC and NDR thank the STScI for DDRF Summer funding to support their research. We thank the referee for a detailed report, which helped us to clarify the discussion of the results of the data analysis in relation to interstellar turbulence. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope which can be found in MAST: 10.17909/26am-1s4. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program 2143.References
@@
- A Full-sky, High-resolution Atlas of Galactic 12 m Dust Emission with WISE. ApJ 781 (1), pp. 5. External Links: Document, 1312.0947, ADS entry Cited by: Appendix B, Appendix C, footnote 14.
@@
- A new method for probing magnetic field strengths from striations in the interstellar medium. MNRAS 481 (4), pp. 5275–5285. External Links: Document, 1810.01559, ADS entry Cited by: §5.3.
- Striations in molecular clouds: streamers or MHD waves?. MNRAS 462 (4), pp. 3602–3615. External Links: Document, 1607.08615, ADS entry Cited by: §5.3.